OpenStudy (anonymous):

sally wants to build a fence around her pool. the pool is 26 feet long by 21 feet wide. the fence is to be 15 feet from the edge of the pool a. What are the outside dimensions of the area surrounding the pool? b. How many feet of fence will sally need? A. 41 ft by 36 ft; 77 ft B. 41 ft by 36 ft; 154 ft C. 56 ft by 51 ft; 214 ft D. 56 ft by 51 ft; 107

OpenStudy (anonymous):

is it B

OpenStudy (anonymous):

Dimensions are 41 feet long and 36 feet wide

OpenStudy (anonymous):

just add 15 feet to both the width, and the length, and then find the perimeter of the fence. length will be (26 + 15 = 41), width will be (21 + 15 = 36). Then use the formula for perimeter (P = 2(l + w) ) to find the perimeter. Your answer choices should already be eliminated down to the first two.

OpenStudy (anonymous):

why does it say im not in this question

OpenStudy (anonymous):

is it B i think its B

OpenStudy (anonymous):

B is correct :)

OpenStudy (anonymous):

yaaaaayyyyyyyy thanxs